OpenStudy (anonymous):

each roller under a conveyor belt has a radius of .5 meters. the roller turns at a rate of 30 revolutions per minute. What is the linear velocity of the conveyor belt.

OpenStudy (anonymous):

The belt can only move as fast as the rollers push it. you have to determine how fast the circumference of the roller is moving. so, determine the circumference (distance in one revolution) and multiply by 30 

OpenStudy (anonymous):

what linear velocity do you need, metre's per minute?

OpenStudy (anonymous):

Meter per second.

OpenStudy (anonymous):

so then it would be 2pi*.5=3.14 and divide that by 30?

OpenStudy (anonymous):

the circumference is right... in one revolution a point on the roller will travel 3.14 metres. now if it undergoes 30 revolutions it won't be a division

OpenStudy (anonymous):

oh I am sorry I meant multiply.

OpenStudy (anonymous):

so it would be 94.2?

OpenStudy (anonymous):

yes, but what units?

OpenStudy (anonymous):

94.2m/minute?

OpenStudy (anonymous):

So then I would need to divide by 60?

OpenStudy (anonymous):

so then it would be 1.57. Thank you for your help! I think I have 2 or 3 more questions and then I am done with all my questions before my test!

OpenStudy (anonymous):

\[\frac{3.14 [metres]}{1 [revolution]}*\frac{30 [revolutions]}{[minute]}\] \[\frac{3.14 [metres]}{1 \cancel {[revolution]}}*\frac{30 \cancel{[revolutions]}}{[minute]}\] \[\frac{94.2 [metres]}{1 [minute]}\]

OpenStudy (anonymous):

\[\frac{94.2 [metres]}{1 \cancel{[minute]}}*\frac{1 \cancel{[minute]}}{60 [seconds]}\]

OpenStudy (anonymous):

Are you learning radians right now?

OpenStudy (anonymous):

Yes. I think so.. I am not very good at trig..